# Estimating non-linear functionals of trawl processes

### v2

@@ 3.1 Limit theorems for $\Psi^{n}(g)$ and $\Lambda^{n}(g)$
    The lower bound in (16), which depends on the memory of $X$, is to prevent that the tail $\int_{N_{n}\Delta_{n}}^{\infty}g(a(s))\mathrm{d}s$ dominates the asymptotics. Thus, unlike Theorem 5 and the first part of Theorem 6, the CLT for $\bar{\Lambda}^{n}(g)$ is no longer robust to long memory.
3.2 A test for $T$-dependent trawls
In this part we briefly exemplify how our results can be used to test for $T$-dependence in a non-parametrical framework. We recall the reader that a stochastic process $(Y_{t})_{t\in\mathbb{R}}$ is said to be $T$-dependent, with $T>0$, if for a given $t_{0}\in\mathbb{R}$, the processes $(Y_{t})_{t\leq t_{0}}$ and $(Y_{t})_{t\geq T+t_{0}}$ are independent (in the sense of finite-dimensional distributions). We have the following characterization of $T$-dependence for trawl processes.
@@ 3.2 A test for $T$-dependent trawls
Since $X$ is an infinitely divisible process, $T$-dependence is equivalent to having that for a given $t_{0}$, $X_{s}$ and $X_{u}$ are independent for all $s\leq t_{0}$ and $u\geq t_{0}+T$ (see for instance Exercise E 12.9 in [16]). Furthermore, by stationarity, the former condition is equivalent to the independence of $X_{t}$ and $X_{0}$ for all $t\geq T$. Now note that, in view of (2) and the independent scatterred property of $L$, the characteristic exponent of the vector $(X_{0},X_{t})$ equals to
@@
In consequence, $X_{t}$ and $X_{0}$ are independent if, and only if,
@@
The preceeding result shows that, under our set-up, $X$ is $T$-dependent if, and only if, $\int_{T}^{\infty}a(s)\mathrm{d}s=0$. Note that this condition is in turn equivalent to $\mathrm{Cov}(X_{T},X_{0})=0$ and thus one can use the sample autocovariance function to test $T$-dependence on $X$. However, it is well known that such a statistic is heavily dependent on the memory of the process. See for instance [6, 7, 10, 21]. To avoid this problem we propose to use another characterization of $T$-dependence, namely
$\int_{T}^{\infty}\rvert a(s)\lvert^{p}\mathrm{d}s=0,\text{ for some (and thus for all) }p>0,$ (18)
which follows from the fact that $a$ is non-negative and non-increasing. Therefore, under the null hypothesis of $T$-dependence we deduce, due to the fact that $\mathrm{Leb}(A)>0$, that
@@
where $g(x)=\lvert x\rvert^{p}$. Based on the previous discussion, we proposed to use the “memory independent” statistic
@@
thanks to (18). This shows the first part of (19). In constrast, if (18) does not hold then necessarily $\int_{0}^{T}\rvert a(s)\lvert^{p}\mathrm{d}s>0$. Theorem 2 now gives that $\tau_{n,p}\overset{\mathbb{P}}{\to}\frac{\int_{T}^{\infty}\rvert a(s)\lvert^{p}\mathrm{d}s}{\int_{0}^{T}\rvert a(s)\lvert^{p}\mathrm{d}s}>0$, from which the second part of (19) follows. ∎
It should be clear from our discussion above and the previous corollary that a large value for $\sqrt{n\Delta_{n}}\tau_{n,p}$ suggests that the data is unlikely to be $T$-dependent.
@@
As pointed out in Remark 2, our proposed statistics for testing (18) is robust to the memory of $X$. Furthermore, as demonstrated in the previous corollary, $\tau_{n,p}$ is more efficient under null. However, it still suffers of some weakness. First, to rigorously use $\tau_{n,p}$ for testing $T$-dependence, a non-trival CLT under the null hipothesis (18) is required. Secondly, it remains to be seen how sensitive is $\tau_{n,p}$ in finite-sample for different values of $p$. These problems requires additional analysis, which we postpone to future research.
4 Proofs
Throughout all our proofs the non-random positive constants will be denoted by the generic symbol $C>0$, and they may change from line to line. We will often write $a\lesssim_{\theta}b$ whenever $a\leq C_{\theta}b$ where the constant $C_{\theta}$ only depends on a given parameter $\theta$. Next, for $i,j=0,1,\ldots$ and $j\geq i$ define
@@ 4 Proofs
Using these sets we introduce the filtrations:
@@
In what is left of this section we set $\mathfrak{t}_{L}=(\gamma,\sigma,\nu)$, the characteristic triplet of $L$, and use the notation $\tau_{n}(s)=[s/\Delta_{n}]\Delta_{n}$ as well as $u_{n}(s)=\left\lceil s/\Delta_{n}\right\rceil\Delta_{n}$, for $s\geq 0.$
4.1 Preliminary estimates and decompositions
In this part we derive some estimates and establish some fundamental decompositions that will be used in a later stage. We also introduce several random variables and processes.
The following estimate for the moments of a homogeneous Lévy basis will be fundamental in our analysis. For a proof see [13] as well as [22] (Corollary 1.2.7).
@@ 4.1 Preliminary estimates and decompositions
Let $L$ be a centered homogeneous Lévy basis with characteristic triplet $\mathfrak{t}_{L}=(\gamma,\sigma,\nu)$ with $\mathbb{E}(\rvert L^{\prime}\rvert^{r})<\infty$ for $r\geq 1$. Then
@@
Next, we observe that the estimation error due to $\hat{a}$ can be represented as
@@
in which $\tilde{X}_{t}=X_{t}-\mathbb{E}(X_{t})$ and $\delta_{k}\tilde{X}=\tilde{X}_{(k+1)\Delta_{n}}-\tilde{X}_{k\Delta_{n}}$. Now, by letting $\tilde{L}(\cdot)=L(\cdot)-\mathbb{E}(L(\cdot)$, define
@@
Based on these random variables and (21) we have the decomposition
@@
Moreover, by using that (once again due to (21))
@@
Without loss of generality we may and do assume that $L$ has mean 0 and so $X$. We will verify that
$E_{n}(l)=(n\Delta_{n})^{r/2}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}[\tilde{a}(l\Delta_{n})]\rvert^{r}),\,\,l=0,\ldots,n-1,$
is bounded by a constant that does not depend neither on $n$ nor on $l$. By (30) and (31) we get that
@@
Note that for every $j\geq l\geq 0$, $(\xi_{j,l,n}^{(1)},\xi_{j,l,n}^{(2)})$, $\rho_{j,l,n}^{(1)}$, and $X_{(j-l)\Delta_{n}}\varrho_{j}$ are martingale difference with respect to $\mathscr{G}_{j}^{n}$, $\mathscr{G}_{j}^{n}\lor\sigma(\beta_{j^{\prime},l^{\prime}}^{(3)})_{0\leq l^{\prime}\leq j^{\prime}\leq j},$ and $\mathscr{F}_{j}^{n}$, respectively. Thus, by Rosenthal’s inequality and Jensen’s inequality we deduce that
@@
$\lesssim_{a}\frac{\mathbb{E}(\rvert X_{0}\rvert^{r})}{(n\Delta_{n})^{r/2}}\sum_{j=l}^{n}\mathbb{E}(\rvert\varrho_{j}\rvert^{r})+\mathbb{E}\left[\frac{1}{n}\sum_{j=l}^{n}\rvert X_{(j-l)\Delta_{n}}\rvert^{r}\right]$
$\lesssim_{r,a,\mathfrak{t}_{L}}\frac{\mathbb{E}(\rvert X_{0}\rvert^{r})}{(n\Delta_{n})^{r/2-1}}+\mathbb{E}(\rvert X_{0}\rvert^{r}),$
where in the last inequality we used (24). Similarly we obtain that
$\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}(\rvert\sum_{j=l}^{n}\rho_{j,l,n}^{(1)}\rvert^{r})$ $\lesssim_{r}\frac{1}{(n\Delta_{n})^{r/2}}\sum_{j=l}^{n}\mathbb{E}(\rvert\rho_{j,l,n}^{(1)}\rvert^{r})+\mathbb{E}\left[\left|\frac{1}{n}\sum_{j=l}^{n}(\beta_{j,l}^{(3)})^{2}\right|^{r/2}\right]$
$\lesssim_{r,a,\mathfrak{t}_{L}}\frac{1}{(n\Delta_{n})^{r/2-1}}+\frac{1}{n}\sum_{j=l}^{n}\mathbb{E}(\rvert\beta_{j,l}^{(3)}\rvert^{r})$
$\lesssim_{r,a,\mathfrak{t}_{L}}\frac{1}{(n\Delta_{n})^{r/2-1}}+1,$
where in the last step we also applied (21). As above we have (once again by Rosenthal’s inequality) for $i=1,2$ that
$E_{n}^{i}(l)\lesssim_{r}\frac{1}{(n\Delta_{n})^{r/2}}\sum_{j=l}^{n}\mathbb{E}(\rvert\xi_{j,n}^{(i)}\rvert^{r})+\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}\left[\left|\sum_{j=l}^{n}\mathbb{E}((\xi_{j,n}^{(i)})^{2}\mid\mathscr{G}_{j-1}^{n})\right|^{r/2}\right].$
Lemma 2 gives us that $\mathbb{E}(\rvert\xi_{j,n}^{(1)}\rvert^{r})\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n}$ and that
$\mathbb{E}(\rvert\xi_{j,n}^{(2)}\rvert^{r})$ $\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n}+\frac{(\Delta_{n}n)^{r/2}}{n}.$
Thus, we are left to show that for $i=1,2$
$E_{n}^{i}(l)^{\prime}=\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}\left[\left|\sum_{j=l}^{n}\mathbb{E}((\xi_{j,n}^{(i)})^{2}\mid\mathscr{G}_{j-1}^{n})\right|^{r/2}\right],$
is bounded uniformly on $n$ and $l$. From Lemma 3 in [17] and (24) we get that
@@
where $\nu_{k}^{(1)}:=\sum_{k^{\prime}=l}^{k-1}\chi_{k^{\prime}}\mathrm{Leb}(\cup_{i=0}^{k^{\prime}-l}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))$. Thus,
$E_{n}^{1}(l)^{\prime}$ $\lesssim_{a,\mathfrak{t}_{L},r}1+\frac{1}{n}\sum_{j=l+1}^{n}\mathbb{E}[\rvert(\beta_{j,j-1,l}^{(2)})^{2}-\mathbb{E}[(\beta_{j,j-1,l}^{(2)})^{2}]\rvert^{r/2}]$
$\lesssim_{a,\mathfrak{t}_{L},r}2,$
and
$E_{n}^{2}(l)^{\prime}\lesssim_{a,\mathfrak{t}_{L},r}$ $1+\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}\left[\left|\sum_{k=l}^{n}\chi_{k}\nu_{k}^{(1)}\right|^{r/2}\right]$
$+\frac{1}{(n\Delta_{n})^{r/2}}\mathbb{E}\left[\left|\sum_{k=l}^{n}[\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})]\mathrm{Leb}(\cup_{i=0}^{k-l}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\right|^{r/2}\right],$
once again by means of (24). If $r>4$ we can apply twice the Rosenthal’s inequality, Jensen’s inequality, and (24) to obtain that
@@
$+\frac{1}{(n\Delta_{n})^{r/2}}\sum_{k=l}^{n}\mathbb{E}(\rvert\chi_{k}\nu_{k}^{(1)}\rvert^{r/2})+\frac{1}{(n\Delta_{n})^{r/4}}\mathbb{E}\left[\left|\frac{1}{n}\sum_{k=l}^{n}(\nu_{k}^{(1)})^{2}\right|^{r/4}\right]$
$\lesssim_{a,\mathfrak{t}_{L},r}$ $1+\frac{1}{(n\Delta_{n})^{r/2-1}}+\frac{1}{(n\Delta_{n})^{r/4}}+\frac{1}{(n\Delta_{n})^{r/2-2}}$
$+\frac{1}{(n\Delta_{n})^{r/4-1}}+\frac{1}{(n\Delta_{n})^{r/4}}\frac{1}{n}\sum_{k=l}^{n}\mathbb{E}(\rvert\nu_{k}^{(1)}\rvert^{r/2})$
$\lesssim_{a,\mathfrak{t}_{L},r}$ $1+\frac{1}{(n\Delta_{n})^{r/2-1}}+\frac{1}{(n\Delta_{n})^{r/4}}+\frac{1}{(n\Delta_{n})^{r/2-2}}+\frac{1}{(n\Delta_{n})^{r/4-1}}.$
Finally, if $r\leq 4$ we can instead use the inequality $\mathbb{E}(\rvert Y\rvert^{r/2})\leq\mathbb{E}(\rvert Y\rvert^{2})^{r/4}$ along with the von Bahr-Esseen inequality to deduce that $E_{n}^{2}(l)^{\prime}$ is bounded up to a constant by
$1+\frac{1}{(n\Delta_{n})^{r/2}}\left|\sum_{k=l}^{n}\mathbb{E}[(\chi_{k}\nu_{k}^{(1)})^{2}]\right|^{r/4}+\frac{1}{(n\Delta_{n})^{r/2}}\sum_{k=l}^{n}\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{r/2}),$
which is in turn uniformly bounded due to (24). ∎
@@
Suppose that $\mathbb{E}(\rvert L^{\prime}\rvert^{2r\lor 4})<\infty$ and that $m_{n}n\rightarrow\jmath\in(0,+\infty].$ Then,
@@
By Lemma 3 we deduce that
$\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{v}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\rvert^{r})\lesssim\frac{1}{(n\Delta_{n})^{r/2}}\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{v}\mathrm{d}s,$
and by the stationarity of $X$ we also obtain that
@@
Furthermore, by Jensen’s inequality and the monoticity of $a$, we deduce that the last summand in (40) is bounded by
$\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{v}[a(\tau_{n}(s))-a(s)]^{r}\mathrm{d}s.$ $\leq a(0)^{v}\int_{0}^{\Delta_{n}}[a(0)-a(s)]\mathrm{d}s$ (41)
$+\int_{0}^{\Delta_{n}}a(s)^{v+1}\mathrm{d}s$
$-\int_{(m_{n}-1)\Delta_{n}}^{m_{n}\Delta_{n}}a(s)^{v+1}\mathrm{d}s$
$\lesssim_{a}\Delta_{n}.$
Next, using that $a$ is integrable and of bounded variation, we can argue as in the proof of Lemma 5.4 in [15] to verify that $xa(x)\rightarrow 0$ as $x\rightarrow\infty$. Consequently
@@
Relation (38) now follows by observing that
@@
thanks to the monoticity of $a$. Finally, if Assumption 1 is satisfied then there is $y_{0}>0$ such that
$\frac{1}{\Delta_{n}}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}[a(l\Delta_{n})-a(s)]\mathrm{d}s$ $\leq\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}\phi(y)\mathrm{d}y$ (42)
$\leq\mathbf{1}_{l\Delta_{n}\leq y_{0}}\Delta_{n}+\mathbf{1}_{l\Delta_{n}>y_{0}}\Delta_{n}(l\Delta_{n})^{-\alpha}.$
@@
4.2 Proof of Theorems 1 - 4
We start by showing the consistency of $\Psi^{n}(g)$. The proof follows the idea of the proof of Theorem 9.4.1 in [11] along with Lemma 3. For the rest of this section we fix $T>0$ and choose $n$ large enough such that $(n-1)\Delta_{n}>T.$

### v3

@@ 3.1. Limit theorems for $\Psi^{n}(g)$ and $\Lambda^{n}(g)$
    The lower bound in (16), which depends on the memory of $X$, prevents the tail $\int_{N_{n}\Delta_{n}}^{\infty}g(a(s))\mathrm{d}s$ from dominating the asymptotics. Thus, unlike Theorem 5 and the first part of Theorem 6, the CLT for $\bar{\Lambda}^{n}(g)$ is no longer robust to long memory.
3.2. A test for $T$-dependent trawls
In this part we briefly illustrate how our results can be used to test for $T$-dependence in a non-parametric framework. We recall that a stochastic process $(Y_{t})_{t\in\mathbb{R}}$ is said to be $T$-dependent, with $T>0$, if for a given $t_{0}\in\mathbb{R}$, the processes $(Y_{t})_{t\leq t_{0}}$ and $(Y_{t})_{t\geq T+t_{0}}$ are independent (in the sense of finite-dimensional distributions). We have the following characterization of $T$-dependence for trawl processes.
@@ 3.2. A test for $T$-dependent trawls
Since $X$ is an infinitely divisible process, $T$-dependence is equivalent to the condition that, for a given $t_{0}$, $X_{s}$ and $X_{u}$ are independent for all $s\leq t_{0}$ and $u\geq t_{0}+T$ (see, for instance, Exercise E 12.9 in [18]). Furthermore, by stationarity, the former condition is equivalent to the independence of $X_{t}$ and $X_{0}$ for all $t\geq T$. Now note that, in view of (2) and the independently scattered property of $L$, the characteristic exponent of the vector $(X_{0},X_{t})$ is equal to
@@
Consequently, $X_{t}$ and $X_{0}$ are independent if, and only if,
@@
The preceding result shows that, under our set-up, $X$ is $T$-dependent if, and only if, $\int_{T}^{\infty}a(s)\mathrm{d}s=0$. Note that this condition is in turn equivalent to $\mathrm{Cov}(X_{T},X_{0})=0$ and thus one can use the sample autocovariance function to test $T$-dependence of $X$. However, it is well known that such a statistic is heavily dependent on the memory of the process. See for instance [6, 7, 11, 23]. To avoid this problem, we propose to use another characterization of $T$-dependence for trawl processes, namely
$\int_{T}^{\infty}\rvert a(s)\lvert^{p}\mathrm{d}s=0,\text{ for some (and hence all) }p>0,$ (18)
which follows from the fact that $a$ is non-negative and non-increasing. Therefore, under the null hypothesis of $T$-dependence, we deduce, since $\mathrm{Leb}(A)>0$, that
@@
where $g(x)=\lvert x\rvert^{p}$. Based on the previous discussion, we propose to use the statistic
@@
thanks to (18). This shows the first part of (19). In contrast, if (18) does not hold then necessarily $\int_{0}^{T}\rvert a(s)\lvert^{p}\mathrm{d}s>0$. Theorem 2 now gives $\tau_{n,p}\overset{\mathbb{P}}{\to}\frac{\int_{T}^{\infty}\rvert a(s)\lvert^{p}\mathrm{d}s}{\int_{0}^{T}\rvert a(s)\lvert^{p}\mathrm{d}s}>0$, from which the second part of (19) follows. ∎
The previous result shows that, unlike the sample autocovariance function, the limiting behaviour of $\tau_{n,p}$ under the alternative of no $T$-dependence is independent of the memory of the process. Consequently, large values of $\sqrt{n\Delta_{n}}\tau_{n,p}$ indicate that the data is unlikely to be $T$-dependent.
@@
As pointed out in Remark 2, our proposed statistic for testing (18) is robust to the memory of $X$. Moreover, in this setting $\tau_{n,p}$ has a smaller estimation error. However, it still suffers from some weaknesses. First, to rigorously use $\tau_{n,p}$ for testing $T$-dependence, a CLT under the null hypothesis (18) is required. Second, it remains to be seen how sensitive $\tau_{n,p}$ is in finite samples for different values of $p$. These issues require additional analysis, which we postpone to future research.
4.  Proofs
The proofs are organised in three parts. We first derive a collection of preliminary estimates and decompositions for the estimator $\hat{a}$, which provide the basic tools used throughout the section. We then use these ingredients to prove the consistency results and to analyse the behaviour of $\Lambda_{t}^{n}(g)$ in the quadratic case. Finally, for our central limit theorems, we use the decompositions obtained in the first part to isolate the leading term, show that all remainder terms are asymptotically negligible, and reduce the problem to a martingale central limit theorem. The technical computations required to approximate the asymptotic variance are deferred to the appendix. As mentioned in the introduction, the appendix may be skipped on a first reading.
Throughout this section, non-random positive constants are denoted by the generic symbol $C>0$, which may change from line to line. We write $a\lesssim_{\theta}b$ whenever $a\leq C_{\theta}b$ where $C_{\theta}$ depends only on the parameter $\theta$. For $i,j=0,1,\ldots$ with $j\geq i$, define
@@ 4.  Proofs
Using these sets, we introduce the filtrations:
@@
In the remainder of this section we set $\mathfrak{t}_{L}=(\gamma,\sigma,\nu)$, the characteristic triplet of $L$, and use the notation $\tau_{n}(s)=[s/\Delta_{n}]\Delta_{n}$ as well as $u_{n}(s)=\left\lceil s/\Delta_{n}\right\rceil\Delta_{n}$, for $s\geq 0.$
4.1. Preliminary estimates and decompositions
In this part we derive some estimates and establish some fundamental decompositions that will be used in a later stage. We also introduce a number of auxiliary random variables and processes.
The following estimate for the moments of a homogeneous Lévy basis will play a key role in our analysis. For a proof see [15] and [24] (Corollary 1.2.7).
@@ 4.1. Preliminary estimates and decompositions
Let $L$ be a centred homogeneous Lévy basis with characteristic triplet $\mathfrak{t}_{L}=(\gamma,\sigma,\nu)$ with $\mathbb{E}(\rvert L^{\prime}\rvert^{r})<\infty$ for $r\geq 1$. Then
@@
Next, we observe that the estimation error $\hat{a}$ admits the decomposition
@@
in which $\tilde{X}_{t}=X_{t}-\mathbb{E}(X_{t})$ and $\delta_{k}\tilde{X}=\tilde{X}_{(k+1)\Delta_{n}}-\tilde{X}_{k\Delta_{n}}$. To analyse this error, we introduce several auxiliary random variables based on the Lévy basis $L$. Let $\tilde{L}(\cdot)=L(\cdot)-\mathbb{E}(L(\cdot)$, define
@@
Using these random variables, we obtain the decomposition
@@
Moreover, by using that
@@
Without loss of generality assume that $L$ (and hence $X$) has mean 0. Define
$E_{n}(l)=(n\Delta_{n})^{r/2}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}[\tilde{a}(l\Delta_{n})]\rvert^{r}),\,\,l=0,\ldots,n-1.$
It suffices to show that $E_{n}(l)$ is bounded uniformly in $n$ and $l$. By (30) and (31) we get that
@@
Note that for every $j\geq l\geq 0$, $(\xi_{j,l,n}^{(1)},\xi_{j,l,n}^{(2)})$, $\rho_{j,l,n}^{(1)}$, and $X_{(j-l)\Delta_{n}}\varrho_{j}$ are martingale differences with respect to $\mathscr{G}_{j}^{n}$, $\mathscr{G}_{j}^{n}\lor\sigma(\beta_{j^{\prime},l^{\prime}}^{(3)})_{0\leq l^{\prime}\leq j^{\prime}\leq j},$ and $\mathscr{F}_{j}^{n}$, respectively. Hence Rosenthal’s inequality (see e.g. Theorem 2.12 in [8]) yields
@@
Using the stationarity of $X$, Jensen’s inequality, and (24) we obtain that
$E_{n}^{4}(l)\lesssim_{r,a,\mathfrak{t}_{L}}\frac{\mathbb{E}(|X_{0}|^{r})}{(n\Delta_{n})^{r/2-1}}+\mathbb{E}(|X_{0}|^{r}).$
Similarly, we deduce that
$E_{n}^{4}(l)\lesssim_{r,a,\mathfrak{t}_{L}}\frac{1}{(n\Delta_{n})^{r/2-1}}+1.$
For $i=1,2$, a further application of Rosenthal’s inequality results in
$E_{n}^{i}(l)\lesssim_{r}\frac{1}{(n\Delta_{n})^{r/2}}\sum_{j=l}^{n}\mathbb{E}(|\xi^{(i)}_{j,n}|^{r})+E_{n}^{i\prime}(l).$
where
$E_{n}^{i}(l)^{\prime}=\mathbb{E}\left[\left|\frac{1}{n\Delta_{n}}\sum_{j=l}^{n}\mathbb{E}((\xi_{j,n}^{(i)})^{2}\mid\mathscr{G}_{j-1}^{n})\right|^{r/2}\right],$
Lemma 2 gives
$\mathbb{E}(\rvert\xi_{j,n}^{(1)}\rvert^{r})\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n},\qquad\mathbb{E}(\rvert\xi_{j,n}^{(2)}\rvert^{r})\lesssim_{r,a,\mathfrak{t}_{L}}\Delta_{n}+\frac{(\Delta_{n}n)^{r/2}}{n}.$
Thus, we are left to bound $E_{n}^{i}(l)^{\prime}$. From Lemma 3 in [19] and (24) we get that
@@
where $\nu_{k}^{(1)}:=\sum_{k^{\prime}=l}^{k-1}\chi_{k^{\prime}}\mathrm{Leb}(\cup_{i=0}^{k^{\prime}-l}\cup_{j=k+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))$. This, together with (24), and Jensen’s inequality shows that $E_{n}^{1}(l)^{\prime}$ is uniformly bounded. Now, if $r>4$ a double application of the Rosenthal’s inequality yields
@@
$+\frac{1}{(n\Delta_{n})^{r/2}}\sum_{k=l}^{n}\mathbb{E}(\rvert\chi_{k}\nu_{k}^{(1)}\rvert^{r/2})+\frac{1}{(n\Delta_{n})^{r/4}}\mathbb{E}\left[\left|\frac{1}{n}\sum_{k=l}^{n}(\nu_{k}^{(1)})^{2}\right|^{r/4}\right].$
Each term on the right-hand side can be shown to be bounded exactly as above. Finally, if $r\leq 4$ we use the inequality $\mathbb{E}(\rvert Y\rvert^{r/2})\leq\mathbb{E}(\rvert Y\rvert^{2})^{r/4}$ along with the von Bahr-Esseen inequality (see [14, 9.3.b]) to deduce that
$E_{n}^{2}(l)^{\prime}\lesssim 1+\frac{1}{(n\Delta_{n})^{r/2}}\left|\sum_{k=l}^{n}\mathbb{E}[(\chi_{k}\nu_{k}^{(1)})^{2}]\right|^{r/4}+\frac{1}{(n\Delta_{n})^{r/2}}\sum_{k=l}^{n}\mathbb{E}(\rvert\chi_{k}^{2}-\mathbb{E}(\chi_{k}^{2})\rvert^{r/2}),$
and the expression above is uniformly bounded due to (24). ∎
@@
Suppose that $\mathbb{E}(\rvert L^{\prime}\rvert^{2r\lor 4})<\infty$ and that $m_{n}\Delta_{n}\rightarrow\jmath\in(0,+\infty].$ Then,
@@
By Lemma 3
$\Delta_{n}\sum_{l=0}^{m_{n}-1}a(l\Delta_{n})^{v}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\rvert^{r})\lesssim\frac{1}{(n\Delta_{n})^{r/2}}\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{v}\mathrm{d}s.$
Using the stationarity of $X$ we also obtain
@@
Furthermore, by Jensen’s inequality and the monotonicity of $a$, we deduce that the last summand in (40) is bounded by
$\int_{0}^{m_{n}\Delta_{n}}a(\tau_{n}(s))^{v}[a(\tau_{n}(s))-a(s)]^{r}\mathrm{d}s\lesssim_{a,r}\Delta_{n}+\int_{0}^{\Delta_{n}}a(s)^{v+1}\mathrm{d}s\lesssim_{a}\Delta_{n}.$ (41)
Moreover, since $a$ is integrable and of bounded variation, it holds that $xa(x)\rightarrow 0$ as $x\rightarrow\infty$ (see the proof of Lemma 5.4 in [17]). Consequently
@@
In view that
@@
relation (38)follows. Finally, if Assumption 1 is satisfied then there is $y_{0}>0$ such that
$\frac{1}{\Delta_{n}}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}[a(l\Delta_{n})-a(s)]\mathrm{d}s\leq\mathbf{1}_{l\Delta_{n}\leq y_{0}}\Delta_{n}+\mathbf{1}_{l\Delta_{n}>y_{0}}\Delta_{n}(l\Delta_{n})^{-\alpha}.$ (42)
@@
4.2. Proof of Theorems 1 - 4
We start by showing the consistency of $\Psi^{n}(g)$. The proof follows the idea of the proof of Theorem 9.4.1 in [12]. For the rest of this section $T>0$ is fixed and choose $n$ large enough such that $(n-1)\Delta_{n}>T.$